The javascript on this page is used to preload the images for links and to swap the images used for links
go to content
layout image go to introduction
go to science datago to mission status reportgo to location / positiongo to 3d data mapsgo to moon VRMLgo to further reading
layout image
Layout imagego to home
go to projectgo to resultsGo to datavisgo to historyGo to scienceGo to educationGo to archivesgo to resources


Lunar Prospector Status Report #52

October 9, 1998 - 1:00 p.m. EST (10:00 a.m. PST)

The Lunar Prospector spacecraft continues to perform very well.

The spacecraft was flipped over this week in two major maneuver intervals. The spacecraft is now flying with its spin axis pointed toward the south ecliptic pole, rather than the north ecliptic pole.

On Oct 5, the spacecraft was reoriented 90 deg to lay in the ecliptic plane, normal to the orbit plane. The attitude was within 0.9 deg of the target attitude and so trim maneuvers were cancelled.

On Oct 7, the spacecraft was reoriented the remaining 90 deg to point towards ecliptic south. The final attitude is within 0.6 deg of the target and again trim maneuvers were cancelled.

Major commanding events are given below.

278/14:11 Pressure Transducer ON

278/14:12 Downlink to OMNI antenna

278/16:05 Start part A of first reor sequence: 255 0.2 sec pulses, A1/A4
jets, precess 54 deg

278/16:44 Start part B of first reor sequence: 174 0.2 sec pulses, A1/A4 jets,
precess 37 deg. Target attitude lat: 0, lon: 282, Actual attd lat: 0.7, lon: 281.5

280/15:17 Start part A of 2nd reor sequence: 255 0.2 sec pulese, A1/A4 jets, precess 53 deg

280/15:49 Start part B of 2nd reor sequence: 179 0.2 sec pulses, A1/A4 jets, precess 37 deg.Target attitude lat: -89.5, lon: 193, Actual lat: -89.1, lon: 156.2

281/15:23 Switch downlink back to Medium Gain Antenna (MGA)

2.5 kg of propellant were used during these maneuvers.

The transmitter continues to be cycled during occultations when the propellant tank heater is on. There were no commanding anomalies that caused any data losses during the past week. There was one instance (DOY 278), when the transmitter could not be commanded off due to a problem at DSS 61, but it did not have an adverse effect on the battery.

On Oct 3, 2 commands were sent to tweak the GRS gain.


Current spacecraft state (0000 GMT 10/09/98):

Orbit: 3292
Downlink: 3600 bps
Spin Rate: 12.08 rpm

Spin Axis Attitude (ecliptic):
Latitude:-89.1 deg
Longitude: 156 deg

Trajectory:
Periapsis Alt: 88.7 km
Apoapsis Alt: 108.3 km

Period: 118 min
Occultations: 41 minutes
Eclipses: 47 minutes
Propellant Remaining: 25.2 kg